112 SMLE HARMONIC MOTION DATA Diameter of sphere 0035m Radius of opl2 Length of 1m 121aam 5 SLSS 2.als 488 1 Percent error Value of g from slope .m Length of peadlm Length of string used Number o ds m 소다 Mass of the spring Force constant of the spring Experine Calculated Prcent ibratios 5 cm 0 cm 5 cm D:3 May someone help me fill in the period and square period for part B and the value of g from slope and percent error for part A?
Answer #1

period = Time / number of oscillations

1.

period = 70.8/50 = 1.416

period2 = 1.4162 = 2.005

2.

period = 70.5/50 = 1.410

period2 = 1.4102 = 1.988

3.

period = 74.5/50 = 1.49

period2 = 1.492 = 2.22